# Revealing Multimodal Causality with Large Language Models

### v1

@@ Abstract
Uncovering cause-and-effect mechanisms from data is fundamental to scientific progress. While large language models (LLMs) show promise for enhancing causal discovery (CD) from unstructured data, their application to the increasingly prevalent multimodal setting remains a critical challenge. Even with the advent of multimodal LLMs (MLLMs), their efficacy in multimodal CD is hindered by two primary limitations: (1) difficulty in exploring intra- and inter-modal interactions for comprehensive causal variable identification; and (2) insufficiency to handle structural ambiguities with purely observational data. To address these challenges, we propose MLLM-CD, a novel framework for multimodal causal discovery from unstructured data. It consists of three key components: (1) a novel contrastive factor discovery module to identify genuine multimodal factors based on the interactions explored from contrastive sample pairs; (2) a statistical causal structure discovery module to infer causal relationships among discovered factors; and (3) an iterative multimodal counterfactual reasoning module to refine the discovery outcomes iteratively by incorporating the world knowledge and reasoning capabilities of MLLMs. Extensive experiments on both synthetic and real-world datasets demonstrate the effectiveness of MLLM-CD in revealing genuine factors and causal relationships among them from multimodal unstructured data. The implementation code and data are available at https://github.com/JinLi-i/MLLM-CD.
@@ 1 Introduction
To the best of our knowledge, COAT [20] is the only existing work that explicitly targets CD from unstructured data using LLMs, marking a significant step forward in this direction. It first extracts causal variables via LLMs and then infers the causal relationships with traditional CD methods. However, it is limited to unimodal data, typically text. In contrast, the rapid proliferation of multimedia content necessitates CD methods capable of effectively handling multimodal information [22, 23] across various domains, such as medical diagnosis [24], where clinical notes, medical images, and lab results collectively inform causal understanding [25, 26, 27]. While it may seem feasible to adapt unimodal approaches like COAT to multimodal scenarios using multimodal LLMs (MLLMs) [28], our findings suggest that such a naïve adaptation falls short. As shown in Figure 2, the adapted COAT struggles with two critical limitations on multimodal unstructured data: (1) it uncovers only a small set of causal factors, and (2) the inferred causal edges remain undirected.
@@
  • We propose MLLM-CD, the first work for causal discovery in multimodal unstructured data with a novel framework that integrates MLLMs into CD pipelines. This significantly extends the scope of CD beyond structured and/or unimodal data settings.
@@
  • We design a novel iterative multimodal counterfactual reasoning module to iteratively refine the inferred causal structures by alleviating structural ambiguities.
@@ 2 Related Work
LLM for Causal Discovery [40, 41, 42, 43, 44, 45, 46] has recently shown great potential in supporting conventional CD pipelines. One of the main approaches is to leverage the general and domain-specific knowledge of LLMs to refine the discovery results [47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57] from statistical CD methods. However, as a post refinement of conventional CD methods, these approaches also suffer from the limitations in unstructured and multimodal data. Motivated by the advanced capabilities of LLMs in understanding natural language descriptions, another line of work focuses on using LLMs to identify causal factors [20] and structures [58, 29, 59, 60] from unstructured data. For instance, Liu et al. [20] introduce a LLM-driven CD framework, which uses LLMs to identify the Markov blanket of the target variable from natural language descriptions. To test the effectiveness of LLMs in causal relationship discovery, Kiciman et al. [29] propose a pairwise prompting method and find that LLMs can be a good candidate for identifying relationships between a pair of variables. Vashishtha et al. [59] further extend this idea to a more accurate and robust triplet-based prompting to resolve the issue of multiple cycles. Nevertheless, these methods are not devised to handle the multimodal nature of data, thus limiting their effectiveness and applicability. Moreover, integrating the contextual understanding and reasoning capabilities of LLMs into statistical CD pipelines—without compromising their statistical rigor—remains a non-trivial challenge. These limitations motivate the development of our proposed approach.
@@ 3.1 Problem Definition
Factor Identification. This task involves identifying the relevant set of latent causal variables $\mathbf{V}$ from multimodal unstructured data $\mathcal{D}$. Crucially, it also requires annotating the value $v_{kj}$ for each variable $V_{j}\in\mathbf{V}$ corresponding to each sample $\mathbf{X}_{k}\in\mathcal{D}$. This process essentially transforms the unstructured multimodal data $\mathcal{D}=\{(\mathbf{x}_{k1},\cdots,\mathbf{x}_{km})\}^{n}_{k=1}$ into a structured dataset $\mathcal{D}_{S}=\{(v_{k1},\cdots,v_{kd})\}^{n}_{k=1}$. Moving beyond solely unimodal factor discovery [20], this step must consider the multimodal nature of the data, and handle: (1) Intra-modal complexity to extract high-level semantic concepts within each modality (e.g., identifying the lesion in a medical image); (2) Inter-modal dependencies to recognize how concepts span or relate across different modalities (e.g., linking the lesion in medical images to the examination findings in clinical notes).
Causal Structure Discovery. Given the identified variables $\mathbf{V}$ and their annotated values in $\mathcal{D}_{S}$, the goal is to infer causal relationships among $\mathbf{V}$ and the target variable $Y$, forming a causal graph $\mathcal{G}=(\mathbf{V}\cup\{Y\},\mathbf{E})$ that best explains the statistical dependencies observed in $\mathcal{D}_{S}$. However, even with the identified variables, this task remains challenging due to issues like Markov equivalence classes [13], where multiple DAGs encode the same conditional independencies.
@@ 3.2 Contrastive Factor Discovery
We propose a CFD module to identify potential causal factors $\mathbf{V}^{(t)}$ and annotate corresponding values in the $t$-th round. This module guides the MLLM $\Psi$ to explore the intra- and inter-modal interactions with contrastive signals, from which it can identify the underlying factors.
@@
Intra-modal Contrastive Exploration: To uncover factors embedded within $i$-th modality, we select the top-$K$ pairs of samples $\mathcal{P}_{i}=\{(\mathbf{x}_{ai},\mathbf{x}_{bi})_{o}\}_{o=1}^{K}$ with the maximum distance $d(a,b)=1-{\rm sim}(\mathbf{e}_{ai},\mathbf{e}_{bi})$, where ${\rm sim}(\mathbf{e}_{ai},\mathbf{e}_{bi})=\mathbf{e}_{ai}\cdot\mathbf{e}_{bi}/(\|\mathbf{e}_{ai}\|\|\mathbf{e}_{bi}\|)$ measures cosine similarity, and $\|\cdot\|$ is the Euclidean norm. These samples are then presented to $\Psi$ with a prompt $\mathbf{p}_{\rm intra}$ (see Appendix D.2) to analyze the underlying intra-model interactions and identify the contributing factors.
@@
Factor Consolidation and Annotation: The contrastive analyses generate multiple sets of candidate factors, which may contain redundancies or overlap with previously identified factors $\mathbf{V}^{(t-1)}$. For deduplication, we prompt $\Psi$ with $\mathbf{p}_{\rm m}$ to merge similar factors and produce a final factor set $\mathbf{V}^{(t)}$. Then, we use the model with prompt $\mathbf{p}_{\rm a}$ to annotate corresponding values, forming a structured dataset $\mathcal{D}_{S}^{(t)}$ for causal structure discovery. The overall process of the CFD module is summarized as:
$\mathbf{V}^{(t)}:=\Psi(\Psi(\{\mathcal{P}_{i}\}^{m}_{i=1},\mathbf{p}_{\rm intra}),\Psi(\mathcal{P}_{\rm x},\mathbf{p}_{\rm inter}),\mathbf{V}^{(t-1)},\mathbf{p}_{\rm m}),\quad\mathcal{D}_{S}^{(t)}:=\Psi(\mathcal{D}^{(t)},\mathbf{V}^{(t)},\mathbf{p}_{\rm a}).$ (1)
@@ 3.3 Causal Structure Discovery
With the identified factors $\mathbf{V}^{(t)}$ and their annotated values in $\mathcal{D}_{S}^{(t)}$, we adopt a statistical CD method $\mathcal{C}$ (e.g., FCI or PC [31]) to infer the causal structure into a causal graph:
$\mathcal{G}^{(t)}=\mathcal{C}(\mathcal{D}_{S}^{(t)},\mathbf{V}^{(t)}\cup\{Y\}).$ (2)
Given the potential presence of unobserved confounders in real-world scenarios, algorithms robust to latent variables, such as FCI algorithm [31] or its variants (e.g., RFCI [66]), are particularly relevant. In this work, we instantiate our framework with the FCI algorithm. However, it can be easily adapted with other algorithms [31] for different causal assumptions.
@@ 3.4 Multimodal Counterfactual Reasoning
The causal graph $\mathcal{G}^{(t)}$ inferred by statistical methods from the derived datasets $\mathcal{D}_{S}^{(t)}$ often contains ambiguity (e.g., represented by undirected or bidirected edges in a partial ancestral graph) due to Markov equivalence classes [13] and potential limitations of finite and noisy observational data. To alleviate this issue, we introduce the MCR module, which leverages the implicit world knowledge and reasoning capabilities of an MLLM $\Psi$ to refine the causal graph iteratively.
The core idea is to use the MLLM to explore “what if” scenarios regarding the uncertainty in $\mathcal{G}^{(t)}$. By simulating interventions on specific factors and predicting how other factors are affected and how the sample changes, we generate counterfactual samples that implicitly encode causal assumptions derived from MLLM’s knowledge. Once validated for plausibility and consistency, they are combined with observational data to provide complementary evidence supporting statistical discovery.
@@
$R_{\rm indep}(\mathbf{S}^{\prime}_{k},V_{a},\mathcal{G}^{(t)})=\frac{\Sigma_{V_{j}\in{\rm NonDesc}(V_{a},\mathcal{G}^{(t)})}\mathbb{I}[\Delta v_{kj}\geq\epsilon]}{|V_{j}\in{\rm NonDesc}(V_{a},\mathcal{G}^{(t)})|}.$ (5)
@@
The validated samples $\mathcal{D}^{(t)}_{CF}$ with $I_{\rm sem}\land I_{\rm causal}=1$ are included for the next round as $\mathcal{D}^{(t+1)}=\mathcal{D}^{(t)}\cup\mathcal{D}^{(t)}_{CF}$. In this way, the observational dataset is effectively augmented using the MLLM’s knowledge and reasoning capabilities, facilitating the iterative refinement of causal factor and structure discovery. The overall process of MLLM-CD is summarized in Algorithm 1 (Appendix C).
@@
4 Experiments
We evaluate both the causal factor and structure discovery performance of MLLM-CD on both synthetic and real-world multimodal datasets, based on state-of-the-art multimodal LLMs including GPT-4o [68], Gemini 2.0 [69], LLaMA 4 Maverick [70], and Grok-2v [71]. Due to the space limit, full comparison results and detailed analyses, including more ablation studies, parameter analysis, time complexity analysis, and case studies, are provided in Appendix D.
@@ 4.1 Experimental Setup
Metrics: Following the common practice in CD [30], we evaluate the performance of both factor identification and causal structure discovery with commonly used metrics. For factor identification, we use: (1) Node Precision (NP), (2) Node Recall (NR), and (3) Node F1-score (NF). For causal structure discovery, we adopt (1) Adjacency Precision (AP), (2) Adjacency Recall (AR), and (3) Adjacency F1-score (AF). To jointly evaluate both perspectives, we use an extended Structural Hamming Distance [80] (ESHD), which incorporates additional penalization for mismatched nodes and edges. The specific definition is offered in Appendix D.3.
@@ 4.2 Analysis with Multimodal Apple Gastronome Dataset
The empirical results of causal discovery on the MAG dataset are shown in Table 1 and Figure 3. The full quantitative results and qualitative comparisons are available in Appendix D.4. MLLM-CD overall outperforms baselines in most cases.

### v2

@@ Abstract
Uncovering cause-and-effect mechanisms from data is fundamental to scientific progress. While large language models (LLMs) show promise for enhancing causal discovery (CD) from unstructured data, their application to the increasingly prevalent multimodal setting remains a critical challenge. Even with the advent of multimodal LLMs (MLLMs), their efficacy in multimodal CD is hindered by two primary limitations: (1) difficulty in exploring intra- and inter-modal interactions for comprehensive causal variable identification; and (2) insufficiency to handle structural ambiguities with purely observational data. To address these challenges, we propose MLLM-CD, a novel framework for multimodal causal discovery from unstructured data. It consists of three key components: (1) a novel contrastive factor discovery module to identify genuine multimodal factors based on the interactions explored from contrastive sample pairs; (2) a statistical causal structure discovery module to infer causal relationships among discovered factors; and (3) an iterative multimodal counterfactual reasoning module to refine the discovery outcomes iteratively by incorporating the world knowledge and reasoning capabilities of MLLMs. Extensive experiments on both synthetic and real-world datasets demonstrate the effectiveness of the proposed MLLM-CD in revealing genuine factors and causal relationships among them from multimodal unstructured data. The implementation code and data are available at https://github.com/JinLi-i/MLLM-CD.
@@ 1 Introduction
To the best of our knowledge, COAT [20] is the only existing work that explicitly targets CD from unstructured data using LLMs, marking a significant step forward in this direction. It first extracts causal variables via LLMs and then infers the causal relationships with traditional CD methods. However, it is limited to unimodal data, typically text. In contrast, the rapid proliferation of multimedia content necessitates CD methods capable of effectively handling multimodal information [22, 23] across various domains, such as medical diagnosis [24], where clinical notes, medical images, and lab results collectively inform causal understanding [25, 26, 27]. While it may seem feasible to adapt unimodal approaches like COAT to multimodal scenarios using multimodal LLMs (MLLMs) [28], our findings suggest that such a naive adaptation falls short. As shown in Figure 2, the adapted COAT struggles with two critical limitations on multimodal unstructured data: (1) it uncovers only a small set of causal factors, and (2) the inferred causal edges remain undirected.
@@
  • We propose MLLM-CD, to the best of our knowledge, the first work for causal discovery in multimodal unstructured data with a novel framework that integrates MLLMs into CD pipelines. This significantly extends the scope of CD beyond structured and/or unimodal data settings for practical use.
@@
  • We design a novel iterative multimodal counterfactual reasoning module to iteratively refine the inferred causal structures by generating multimodal counterfactuals.
@@ 2 Related Work
LLM for Causal Discovery [40, 41, 42, 43, 44, 45, 46, 47] has recently shown great potential in supporting conventional CD pipelines. One of the main approaches is to leverage the general and domain-specific knowledge of LLMs to refine the discovery results [48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58] from statistical CD methods. However, as a post-refinement of conventional CD methods, these approaches also suffer from limitations in unstructured and multimodal data. Motivated by the advanced capabilities of LLMs in understanding natural language descriptions, another line of work focuses on using LLMs to identify causal factors [20] and structures [59, 29, 60, 61] from unstructured data. For instance, Liu et al. [20] introduce a LLM-driven CD framework, which uses LLMs to identify the Markov blanket of the target variable from natural language descriptions. To test the effectiveness of LLMs in causal relationship discovery, Kiciman et al. [29] propose a pairwise prompting method and find that LLMs can be a good candidate for identifying relationships between a pair of variables. Vashishtha et al. [60] further extend this idea to a more accurate and robust triplet-based prompting to resolve the issue of multiple cycles. Nevertheless, these methods are not devised to handle the multimodal nature of data, thus limiting their effectiveness and applicability. Moreover, integrating the contextual understanding and reasoning capabilities of LLMs into statistical CD pipelines—without compromising their statistical rigor—remains a non-trivial challenge. These limitations motivate the development of our proposed approach.
Causal Representation Learning (CRL) [62, 63, 64] aims to extract high-level representations and causal dependencies from low-level observational data, such as images or videos. Much of the work in this area is related to disentangled representation learning [65, 66, 67, 68], which tries to map distinct factors of variation in the data to separate latent variables. In addition to observational data, using data generated under interventions or multiple environments has also been explored to learn more robust and causally meaningful representations [69, 70, 71]. While CRL has shown promise in providing provable identifiability of latent causal variables under certain assumptions, it remains a developing field with open challenges [72, 69, 73] in practical use. In this paper, we aim to identify human-interpretable causal variables and corresponding causal structures from multimodal unstructured data by leveraging the contextual understanding and reasoning capabilities of MLLMs.
@@ 3.1 Problem Definition
Factor Identification. This task involves identifying the relevant set of causal variables $\mathbf{V}$ from multimodal unstructured data $\mathcal{D}$. Crucially, it also requires annotating the value $v_{kj}$ for each variable $V_{j}\in\mathbf{V}$ corresponding to each sample $\mathbf{X}_{k}\in\mathcal{D}$. This process essentially transforms the unstructured multimodal data $\mathcal{D}=\{(\mathbf{x}_{k1},\cdots,\mathbf{x}_{km})\}^{n}_{k=1}$ into a structured dataset $\mathcal{D}_{\rm S}=\{(v_{k1},\cdots,v_{kd})\}^{n}_{k=1}$. Moving beyond solely unimodal factor discovery [20], this step must consider the multimodal nature of the data, and handle: (1) Intra-modal complexity to extract high-level semantic concepts within each modality (e.g., identifying the lesion in a medical image); (2) Inter-modal dependencies to recognize how concepts span or relate across different modalities (e.g., linking the lesion in medical images to the examination findings in clinical notes).
Causal Structure Discovery. Given the identified variables $\mathbf{V}$ and their annotated values in $\mathcal{D}_{\rm S}$, the goal is to infer causal relationships among $\mathbf{V}$ and the target variable $Y$, forming a causal graph $\mathcal{G}=(\mathbf{V}\cup\{Y\},\mathbf{E})$ that best explains the statistical dependencies observed in $\mathcal{D}_{\rm S}$. However, even with the identified variables, this task remains challenging due to issues like Markov equivalence classes [13], where multiple DAGs encode the same conditional independencies.
@@ 3.2 Contrastive Factor Discovery
We propose a CFD module to identify potential causal factors $\mathbf{V}^{(t)}$ and annotate corresponding values in the $t$-th round. This module adopts an MLLM, denoted as $\Psi$, to explore the intra- and inter-modal interactions with contrastive signals, from which it can identify the underlying factors.
@@
Intra-modal Contrastive Exploration: To uncover factors embedded within $i$-th modality, we select the top-$K$ pairs of samples $\mathcal{P}_{i}=\{(\mathbf{x}_{ai},\mathbf{x}_{bi})_{o}\}_{o=1}^{K}$ with the maximum distance $d(a,b)=1-{\rm sim}(\mathbf{e}_{ai},\mathbf{e}_{bi})$, where ${\rm sim}(\mathbf{e}_{ai},\mathbf{e}_{bi})=\mathbf{e}_{ai}\cdot\mathbf{e}_{bi}/(\|\mathbf{e}_{ai}\|\|\mathbf{e}_{bi}\|)$ measures cosine similarity, and $\|\cdot\|$ is the Euclidean norm. These samples are then presented to $\Psi$ with a prompt $\mathbf{p}_{\rm intra}$ (see Appendix D.2) to analyze the underlying intra-modal interactions and identify the contributing factors.
@@
Factor Consolidation and Annotation: The contrastive analyses generate multiple sets of candidate factors, which may contain redundancies or overlap with previously identified factors $\mathbf{V}^{(t-1)}$. For deduplication, we prompt $\Psi$ with $\mathbf{p}_{\rm m}$ to merge similar factors and produce a final factor set $\mathbf{V}^{(t)}$. Then, we use the model with prompt $\mathbf{p}_{\rm a}$ to annotate corresponding values, forming a structured dataset $\mathcal{D}_{\rm S}^{(t)}$ for causal structure discovery. The overall process of the CFD module is summarized as:
$\mathbf{V}^{(t)}:=\Psi(\Psi(\{\mathcal{P}_{i}\}^{m}_{i=1},\mathbf{p}_{\rm intra}),\Psi(\mathcal{P}_{\rm x},\mathbf{p}_{\rm inter}),\mathbf{V}^{(t-1)},\mathbf{p}_{\rm m}),\quad\mathcal{D}_{\rm S}^{(t)}:=\Psi(\mathcal{D}^{(t)},\mathbf{V}^{(t)},\mathbf{p}_{\rm a}).$ (1)
@@ 3.3 Causal Structure Discovery
With the identified factors $\mathbf{V}^{(t)}$ and their annotated values in $\mathcal{D}_{\rm S}^{(t)}$, we adopt a statistical CD method $\mathcal{C}$ (e.g., FCI or PC [31]) to infer the causal structure into a causal graph:
$\mathcal{G}^{(t)}=\mathcal{C}(\mathcal{D}_{\rm S}^{(t)},\mathbf{V}^{(t)}\cup\{Y\}).$ (2)
Given the potential presence of unobserved confounders in real-world scenarios, algorithms robust to latent variables, such as FCI algorithm [31] or its variants (e.g., RFCI [79]), are particularly relevant. In this work, we instantiate our framework with the standard FCI algorithm. However, it can be easily adapted with other algorithms [31] for different causal assumptions and data types [80].
@@ 3.4 Multimodal Counterfactual Reasoning
The causal graph $\mathcal{G}^{(t)}$ inferred by statistical methods from the derived datasets $\mathcal{D}_{\rm S}^{(t)}$ often contains ambiguity (e.g., represented by undirected edges in a partial ancestral graph) due to potential limitations of finite and noisy observational data. To alleviate this issue, we introduce the MCR module, which leverages the implicit world knowledge and reasoning capabilities of an MLLM $\Psi$ to refine the causal graph iteratively.
The core idea is to use the MLLM to explore “what if” scenarios regarding the uncertainty in $\mathcal{G}^{(t)}$, thereby generating counterfactuals to support genuine causal relationships. By simulating interventions on specific factors and predicting how other factors are affected and how the sample changes, we generate counterfactual samples that implicitly encode causal assumptions derived from MLLM’s knowledge. Once validated for plausibility and consistency, they are combined with observational data to provide complementary evidence supporting statistical discovery.
@@
$R_{\rm indep}(\mathbf{S}^{\prime}_{k},V_{a},\mathcal{G}^{(t)})=\frac{\Sigma_{V_{j}\in{\rm NonDesc}(V_{a},\mathcal{G}^{(t)})}\mathbb{I}[\Delta v_{kj}\geq\epsilon]}{|{\rm NonDesc}(V_{a},\mathcal{G}^{(t)})|}.$ (5)
@@
Note that, since the current graph $\mathcal{G}^{(t)}$ used for the check is not ground truth and may contain errors, a non-zero threshold $\tau_{\rm causal}$ is necessary. This threshold serves as a trade-off between injecting MLLM’s knowledge and maintaining consistency with a potentially improvable graph. The validated samples $\mathcal{D}^{(t)}_{\rm CF}$ with $I_{\rm sem}\land I_{\rm causal}=1$ are included for the next round as $\mathcal{D}^{(t+1)}=\mathcal{D}^{(t)}\cup\mathcal{D}^{(t)}_{\rm CF}$. In this way, the observational dataset is effectively augmented using the MLLM’s knowledge and reasoning capabilities, facilitating the iterative refinement of causal factor and structure discovery. The overall process of MLLM-CD is summarized in Algorithm 1 (Appendix C).
Despite the promise of modern MLLMs in demonstrating high-level general knowledge and reasoning capabilities [15], it is important to acknowledge that they may still fall short in challenging scenarios requiring cross-domain or domain-specific expertise. In such cases, we recommend applying stricter validation to filter out noisy counterfactuals and integrating with strategies like retrieval-augmented generation (RAG) [82, 83] or knowledge graph (KG) grounding [84, 85] to enhance the model’s knowledge base and reasoning capabilities.
4 Experiments
We evaluate both the causal factor and structure discovery performance of MLLM-CD on both synthetic and real-world multimodal datasets, based on state-of-the-art multimodal LLMs including GPT-4o [86], Gemini 2.0 [87], LLaMA 4 Maverick [88], and Grok-2v [89]. Due to the space limit, full comparison results and detailed analyses, including more ablation studies, parameter analysis, time complexity and scalability analysis, and case studies, are provided in Appendix D.
@@ 4 Experiments
Pairwise - - - $0.47\pm 0.12$ $0.38\pm 0.13$ $0.40\pm 0.09$ $34.08\pm 8.76$
Triplet - - - $0.43\pm 0.20$ $0.40\pm 0.14$ $0.39\pm 0.07$ $39.58\pm 15.00$
@@ 4.1 Experimental Setup
Metrics: Following the common practice in CD [30], we evaluate the performance of both factor identification and causal structure discovery with commonly used metrics. For factor identification, we use: (1) Node Precision (NP), (2) Node Recall (NR), and (3) Node F1-score (NF). For causal structure discovery, we adopt (1) Adjacency Precision (AP), (2) Adjacency Recall (AR), and (3) Adjacency F1-score (AF). To jointly evaluate both perspectives, we use an extended Structural Hamming Distance [98] (ESHD), which incorporates additional penalization for mismatched nodes and edges. More details are offered in Appendix D.3.
@@ 4.2 Analysis with Multimodal Apple Gastronome Dataset
The empirical results of causal discovery on the MAG dataset are shown in Table 1 and Figure 3. The full comparison results are available in Appendix D.4. MLLM-CD overall outperforms baselines in most cases.